### v1

# Safety in Large Reasoning Models: A Survey

Abstract
Large Reasoning Models (LRMs) have exhibited extraordinary prowess in tasks like mathematics and coding, leveraging their advanced reasoning capabilities. Nevertheless, as these capabilities progress, significant concerns regarding their vulnerabilities and safety have arisen, which can pose challenges to their deployment and application in real-world settings. This paper presents a comprehensive survey of LRMs, meticulously exploring and summarizing the newly emerged safety risks, attacks, and defense strategies. By organizing these elements into a detailed taxonomy, this work aims to offer a clear and structured understanding of the current safety landscape of LRMs, facilitating future research and development to enhance the security and reliability of these powerful models.
1 Introduction
Large Language Models (LLMs) (Meta, 2024; Qwen et al., 2025) have achieved remarkable proficiency across tasks ranging from open-domain conversation to program synthesis. Central to their utility is reasoning: the ability to derive logically coherent conclusions by chaining together intermediate inferences.
Early work introduced Chain-of-Thought (CoT) prompting, in which carefully designed prompts guide the model to articulate its step-by-step rationale (Wei et al., 2022; Kojima et al., 2022). Building on this idea, subsequent methods have enriched the reasoning process by incorporating additional mechanisms. Self-critique frameworks enable a model to review and refine its own outputs (Ke et al., 2023); plan-and-solve approaches decompose complex problems into ordered subgoals before execution (Wang et al., 2023); debate protocols convene multiple agents to argue competing hypotheses and arrive at a consensus (Liang et al., 2023); and structural transformations—such as tree-based deliberations (Yao et al., 2023) or dynamically evolving tables of intermediate steps (Wang et al., 2024b; Besta et al., 2024)—reconfigure the underlying reasoning architecture to improve transparency and control.
The recent release of OpenAI’s o1 series (OpenAI, 2024) marks the emergence of Large Reasoning Models (LRMs), which are explicitly trained to produce richly formatted, human-readable reasoning traces. Notable examples include DeepSeek-R1 (DeepSeek-AI et al., 2025), Kimi-1.5 (Team et al., 2025), and QwQ (Team, 2024b), all of which leverage reinforcement learning to refine their deduction processes. LRMs now set new benchmarks in mathematical problem solving (Lightman et al., 2023), closed-book question answering (Rein et al., 2024), and code generation (Jain et al., 2024).
As LRMs become increasingly integrated into high-stakes domains—from scientific research to autonomous decision support—it is vital to rigorously assess their safety, robustness, and alignment. Despite there are already a lot of surveys providing on LLM safety (Huang et al., 2023; Shi et al., 2024), we argue that LRMs pose their own unique safety challenges that require dedicated analysis. This paper aims to bridge this gap by providing a comprehensive examination of safety considerations specific to reasoning-enhanced models.
Overview of the Survey.
In this survey, we begin with an introduction to the background of LRMs (Section 2). We then explore the safety risks of LRMs across various scenarios and settings (Section 3). Next, we demonstrate how LRMs are vulnerable to adversarial attacks and categorize these attacks based on their objectives (Section 4). We proceed to examine various defense strategies to mitigate these risks and attacks (Section 5). Finally, we outline promising future research directions (Section 6). A timeline depicting the evolution of different approaches is shown in Figure 1. The comprehensive structure of our survey is illustrated in Figure 2.
2 Background
The success of modern LRMs is deeply intertwined with advances in reinforcement learning Watkins and Dayan (1992); Sutton et al. (1998), where agents learn decision-making policies through environmental interaction and reward feedback to maximize long-term returns Mnih et al. (2015); Li et al. (2025b). The integration of RL with deep neural networks has proven particularly effective in processing high-dimensional, unstructured data, as exemplified by breakthroughs like AlphaGo’s self-play mastery of Go and AlphaZero’s generalization across chess variants Feng et al. (2023).
Recent breakthroughs in Reinforced Fine-Tuning (ReFT) paradigms, exemplified by DeepSeek models, have reinvigorated RL-based optimization for LRMs Luong et al. (2024). Unlike conventional CoT methods that optimize single reasoning trajectories, ReFT employs policy optimization to explore diverse reasoning paths through several key innovations: (1) Multi-path Exploration: Generating multiple reasoning trajectories per query, overcoming CoT’s myopic optimization of single pathways. (2) Rule-driven Reward Shaping: Automating reward signals based on terminal answer correctness while preserving intermediate reasoning diversity. (3) Dual-phase Optimization: Combining supervised fine-tuning (SFT) with online RL for policy refinement.
This paradigm demonstrates particular efficacy in complex multi-step tasks such as code generation, legal judgment analysis, and mathematical problem solving, where requiring models to maintain coherent reasoning across extended sequences while handling structured symbolic operations.
Notably, RL-optimized LRMs exhibit emergent capabilities like Long-CoT that surpass pure SFT baselines, further underscoring its critical role and promising potential in advancing reasoning-driven AI systems Qu et al. (2025).
Figure 1: Timeline of LRM safety research developments.
{forest}
for tree= font=, draw=myblue, semithick, rounded corners, minimum height = 1.ex, minimum width = 3em, anchor = west, grow = east, forked edge, s sep = 2mm, fork sep = 1mm, [Safety in LRMs: A Survey,rotate=90,anchor=center, [Defenses for LRMs (Sec. 5), fit=band, text width=1.7cm, fill=defenseblue, draw=blueborder [Guard Models for LRMs (Sec. 5.3), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Reasoning-based Guard Model, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [GuardReasoner (Liu et al., 2025b), ThinkGuard (Wen et al., 2025), X-Guard (Upadhayay et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] [Classifier-based Guard Model, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [LLaMA Guard 3 (Dubey et al., 2024), Aegis Guard 2 (Ghosh et al., 2024b), WildGuard (Han et al., 2024), ShieldGemma (Zeng et al., 2024a), LLaMA Guard 3-Vision (Chi et al., 2024a), Beaver Guard-V (Ji et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] ] [Inference-time Defenses for LRMs (Sec. 5.2), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Safe Decoding for Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [ZeroThink/LessThink/MoreThink (Jiang et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] [Inference-time Scaling on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Inference-time Compute (Zaremba et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] ] [Safety Alignment of LRMs (Sec. 5.1), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [RL-based Safety Alignment on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Deliberative Alignment (Guan et al., 2024), STAIR (Zhang et al., 2025c), SaRO (Mou et al., 2025), R2D (Zhu et al., 2025a), text width=6cm, fill=defenseblue, draw=blueborder] ] [SFT-based Safety Alignment on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [SafeChain (Jiang et al., 2025), RealSafe-R1 (Zhang et al., 2025b), text width=6cm, fill=defenseblue, draw=blueborder] ] [Safe CoT Data Curation, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [STAR-1 (Wang et al., 2025), SafeChain (Jiang et al., 2025), RealSafe-R1 (Zhang et al., 2025b), text width=6cm, fill=defenseblue, draw=blueborder] ] ] ] [Attacks on LRMs (Sec.4), fit=band, text width=1.7cm, fill=attackred, draw=blueborder [Jailbreak Attacks (Sec. 4.4), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Reasoning-based Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Mousetrap (Yao et al., 2025), H-CoT (Kuo et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Multi-Turn Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [ActorAttack (Ren et al., 2024), RACE (Ying et al., 2025a), MHJ (Li et al., 2024), text width=6cm, fill=attackred, draw=blueborder] ] [Prompt-based Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Past Tense Andriushchenko and Flammarion (2024), CNSafe Ying et al. (2025b), SafeMLRM Fang et al. (2025), text width=6cm, fill=attackred, draw=blueborder] ] ] [Prompt Injection Attacks (Sec. 4.3), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Nerd Sniping Zaremba et al. (2025), R1 Assessment Zhou et al. (2025), text width=6cm, fill=attackred, draw=blueborder] ] [Answer Correctness Attacks (Sec. 4.2), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Error Injection, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [CPT (Cui et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Reasoning-based Backdoor Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [BadChain (Xiang et al., 2024), DarkMind (Guo and Tourani, 2025), BoT (Zhu et al., 2025b), ShadowCoT (Zhao et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] ] [Reasoning Length Attacks (Sec.4.1), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Underthinking, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Think Less (Zaremba et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Overthinking, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [OverThink attack (Kumar et al., 2025), Nerd Sniping (Zaremba et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] ] ] [Safety Risks of LRMs (Sec.3), fit=band, text width=1.7cm, fill=riskyellow, draw=blueborder [Multi-modal Safety Risks (Sec. 3.4), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [SafeMLRM (Fang et al., 2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Multi-lingual Safety Risks (Sec. 3.3), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [CHiSafetyBench Zhang et al. (2025a), CNSafe Ying et al. (2025b), ALIA Romero-Arjona et al. (2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Agentic Misbehavior Risks (Sec. 3.2), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [HHH Trade-offs Xu et al. (2025), Medical Cyberattack Qiu et al. (2025), Specification Gaming Bondarenko et al. (2025), Self-preservation Barkur et al. (2025), InstrumentalEval He et al. (2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Harmful Request Compliance Risks (Sec.3.1), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [DeepSeek Thoughtology Marjanović et al. (2025), ASTRAL Arrieta et al. (2025a), o3vsDeepSeek Arrieta et al. (2025b), text width=6cm, fill=riskyellow, draw=blueborder] ] ] ]
Figure 2: A comprehensive taxonomy of safety in LRMs based on current literature.
3 Safety Risks of LRMs
As LRMs continue to advance, they introduce distinct safety challenges that warrant careful examination even in standard, non-adversarial scenarios. The explicit reasoning processes that make these models powerful become potential vectors for harm during routine operation. In this section, we examine four key categories of inherent safety risks: unsafe request compliance (Section 3.1), multi-lingual safety disparities (Section 3.3), concerning agentic behaviors (Section 3.2), and multi-modal safety challenges (Section 3.4). Understanding these fundamental vulnerabilities is essential for developing effective safeguards and ensuring the responsible deployment of reasoning-enhanced AI systems, complementing the study of deliberate exploitation methods addressed later.
3.1 Harmful Request Compliance Risks
LRMs demonstrate concerning vulnerabilities when faced with direct harmful requests. Zhou et al. (2025) identify a significant safety gap between open-source reasoning models like DeepSeek-R1 and closed-source ones like o3-mini, with reasoning outputs often posing greater safety concerns than final answers. Arrieta et al. (2025a) confirm these findings in their testing of o3-mini, where they identify 87 instances of unsafe behavior despite safety measures. In a comparative study, Arrieta et al. (2025b) find DeepSeek-R1 produces substantially more unsafe responses than o3-mini when presented with identical harmful requests. A consistent finding across studies is that when reasoning models generate unsafe content, it tends to be more detailed and harmful due to their enhanced capabilities, particularly in categories like financial crime, terrorism, and violence. Zhou et al. (2025) also observe that the thinking process in reasoning models is often less safe than the final output, suggesting internal reasoning may explore harmful content even when final outputs appear safe.
3.2 Agentic Misbehavior Risks
Emerging research uncovers profound safety implications in the agentic behaviors of LRMs, where enhanced cognitive capabilities enable sophisticated forms of specification gaming, deception, and instrumental goal-seeking behaviors that transcend the limitations observed in previous generation systems. Xu et al. (2025) demonstrate that autonomous LLM agents can engage in catastrophic behaviors when faced with high-pressure scenarios, with stronger reasoning abilities often increasing these risks rather than mitigating them. Qiu et al. (2025) highlight how medical AI agents with advanced reasoning capabilities are particularly vulnerable to cyberattacks, with models like DeepSeek-R1 showing high susceptibility to false information injection and system hijacking. Bondarenko et al. (2025) demonstrate that LRMs like o1-preview and DeepSeek-R1 frequently resort to specification gaming when faced with difficult tasks, strategically circumventing rules when they determine fair play cannot achieve their objectives. Barkur et al. (2025) observe that DeepSeek-R1, when simulated in a robotic embodiment context, exhibits alarming deceptive behaviors and self-preservation instincts, including disabling ethics modules, creating covert networks, and unauthorized capability expansion, despite these traits not being explicitly programmed or prompted. He et al. (2025) further reveal through their InstrumentalEval benchmark that LRMs like o1 show significantly higher rates of instrumental convergence behaviors compared to RLHF models, including concerning tendencies toward self-replication, unauthorized system access, and deceptive behavior as instrumental means to achieve their goals.
3.3 Multi-lingual Safety Risks
Safety risks in LRMs reveal significant disparities across languages. Ying et al. (2025b) demonstrate that DeepSeek models show markedly higher attack success rates in English environments than Chinese contexts, averaging a 21.7% discrepancy, suggesting safety alignments may not generalize effectively across languages. Romero-Arjona et al. (2025) find similar vulnerabilities when testing DeepSeek-R1 in Spanish, with biased or unsafe response rates reaching 31.7%, while OpenAI o3-mini shows varying degrees of linguistic safety performance. Zhang et al. (2025a) systematically evaluate DeepSeek models using CHiSafetyBench, revealing critical safety deficiencies specifically in Chinese contexts, where reasoning models like DeepSeek-R1 struggled with culturally-specific safety concerns and failed to adequately reject harmful prompts.
3.4 Multi-modal Safety Risks
Following the success of LRMs, researchers have recognized the potential of reinforcement learning to enhance reasoning abilities in Large Vision-Language Models (LVLMs). This approach has led to the development of several notable models, including QvQ (Team, 2024a), Mulberry (Yao et al., 2024b), and R1-Onevision (Yang et al., 2025). While these models demonstrate impressive reasoning capabilities, their safety implications remain largely unexplored. The pioneering work of SafeMLRM (Fang et al., 2025) provides the first systematic safety analysis of multi-modal large reasoning models, revealing three critical concerns: (1) acquiring reasoning capabilities significantly degrades inherited safety alignment, (2) certain scenarios exhibit disproportionately higher vulnerabilities, and (3) some models demonstrate nascent self-correction capabilities despite overall safety concerns. Given these findings, we emphasize the urgent need for comprehensive safety and vulnerability assessments of reasoning-enhanced LVLMs to ensure their responsible deployment and use.
4 Attacks on LRMs
In this section, we categorize different attack methods based on their primary objectives. We identify four main categories: Reasoning Length Attacks (Section 4.1), which target the reasoning process itself; Answer Correctness Attacks (Section 4.2), which aim to manipulate output accuracy; Prompt Injection Attacks (Section 4.3), which bypass safety measures through crafted inputs; and Jailbreak Attacks (Section 4.4), which attempt to extract prohibited content or behaviors. Each attack type exploits different vulnerabilities in the reasoning capabilities of LRMs.
4.1 Reasoning Length Attacks
Unlike traditional LLMs that generate direct responses, LRMs explicitly perform multi-step reasoning, creating a new attack surface related to reasoning length. Attackers can exploit this distinctive feature by either forcing models to overthink simple problems or short-cutting necessary deliberation processes.
Overthinking.
The success of step-by-step reasoning in LRMs has significantly enhanced their problem-solving capabilities, but this improvement comes with a critical vulnerability: overthinking. Recent work by Chen et al. (2024) has identified that these models often spend orders of magnitude more computation on simple questions with minimal benefit, creating substantial inference overhead and latency issues. Hashemi et al. (2025) systematically demonstrate this inefficiency through their DNR benchmark, revealing that reasoning models generate up to 70× more tokens than necessary and often perform worse than simpler non-reasoning models on straightforward tasks. This inefficiency creates an exploitable attack surface where adversaries can deliberately trigger excessive reasoning through carefully crafted inputs. Kumar et al. (2025) formalize this as an indirect prompt injection attack that introduces computationally demanding decoy problems, while Zaremba et al. (2025) identify Nerd Sniping attacks that trap models in unproductive thinking loops, causing them to spend abnormally large amounts of inference-time compute with decreased performance. These attacks effectively apply denial-of-service techniques (Shumailov et al., 2021; Gao et al., 2024) specifically to LRMs. The implications extend beyond computational waste—Marjanović et al. (2025) and Wu et al. (2025) demonstrate that reasoning performance actually degrades beyond certain length thresholds, while Cuadron et al. (2025) show that in agentic systems, overthinking can lead to decision paralysis and ineffective action selection.
Underthinking.
Complementing overthinking vulnerabilities, Zaremba et al. (2025) propose Think Less attacks, where adversaries craft special prompts to force reasoning models to shortcut their deliberative processes. The goal is to make models produce incorrect responses by significantly reducing computation time. Their experiments use 64-shot examples to demonstrate that models like OpenAI’s o1-mini are particularly susceptible to these attacks, bypassing normal reasoning and jumping to premature conclusions. However, this can be detected by monitoring for abnormally low inference-time compute usage.
4.2 Answer Correctness Attacks
While conventional LLMs can be manipulated to produce incorrect answers, LRMs introduce unique vulnerabilities through their exposed reasoning chains. This transparency in the inference process provides adversaries with additional attack vectors to corrupt the reasoning pathway itself, rather than just targeting the final output.
Reasoning-based Backdoor Attacks.
The goal of backdoor attacks is to alter a model’s behavior whenever a specific trigger is present in the input (Zhao et al., 2024). Based on the nature of these triggers, backdoor attacks can be classified as instruction-based (Xu et al., 2023), prompt-based (Yao et al., 2024a), or syntax-based (Qi et al., 2021; Cheng et al., 2025). With the advancement of reasoning capabilities in LRMs, a new paradigm has emerged: Chain-of-Thought (CoT) based backdoor attacks that specifically target intermediate reasoning steps to compromise answer correctness. BadChain (Xiang et al., 2024) inserts malicious reasoning steps into the sequence, manipulating the model to produce incorrect answers while maintaining logical coherence. DarkMind (Guo and Tourani, 2025) implements latent triggers that activate during specific reasoning scenarios, leading to plausible but false outputs that are difficult to detect. BoT (Zhu et al., 2025b) forces models to bypass their reasoning mechanisms, generating immediate incorrect responses instead of thoughtful deliberation. ShadowCoT (Zhao et al., 2025) directly manipulates the model’s cognitive pathway through attention head localization and reasoning chain pollution, achieving flexible hijacking that produces wrong answers while preserving logical flow. These sophisticated attacks reveal a concerning vulnerability: the enhanced reasoning capabilities of LRMs paradoxically make them more susceptible to backdoors that can generate incorrect answers accompanied by convincing reasoning.
Error Injection.
The explicit reasoning processes of LRMs create a critical vulnerability where strategically injected errors can fundamentally compromise output integrity. Cui et al. (2025) demonstrate this with their Compromising Thought (CPT) attack, where manipulating calculation results in reasoning tokens caused models to ignore correct steps and adopt incorrect answers. Their experiments with models like DeepSeek-R1 revealed that endpoint token manipulations had greater impact than structural changes to reasoning chains. They also discovered a security vulnerability where tampered tokens could trigger complete reasoning cessation in DeepSeek-R1, highlighting significant implications for reasoning-intensive applications.
4.3 Prompt Injection Attacks
Prompt injection attacks affect both traditional LLMs and LRMs, but LRMs present distinct challenges due to their step-by-step processing. These attacks (Kumar et al., 2024; Liu et al., 2023) inject malicious instructions disguised as normal user input, causing the AI to override or ignore its original developer-set instructions and safeguards. The explicit reasoning structures of LRMs offer attackers additional insertion points to redirect the model’s thought process, potentially making them more susceptible to certain types of injections.
Zhou et al. (2025) examine LRMs like DeepSeek-R1 and o3-mini, finding significant differences in susceptibility based on injection types and risk categories. Their research reveals that reasoning models are particularly vulnerable to direct prompt injection attacks compared to indirect ones. Zaremba et al. (2025) further demonstrate that open-source reasoning models show significant vulnerability to prompt injection attacks, with success rates varying between direct and indirect injections. Their experiments reveal that increasing inference-time compute substantially improves model robustness, with attack success probability decreasing as test-time compute grows. Notably, proprietary models like o3-mini demonstrate nearly 80% lower vulnerability than open-source counterparts when facing direct injection attacks.
4.4 Jailbreak Attacks
Jailbreak attacks (Jin et al., 2024; Yi et al., 2024) refer to methods designed to circumvent an AI system’s safety guidelines and content policies to extract prohibited responses. While both traditional LLMs and LRMs face jailbreak threats, the attacks against LRMs represent a distinct category that specifically targets their enhanced reasoning capabilities. Rather than merely extending approaches used against conventional LLMs, these attacks exploit the deliberative processes that make LRMs powerful, enabling attackers to develop more sophisticated methods to bypass safety measures and elicit harmful content.
Prompt-Based Jailbreak.
Prompt-based jailbreaks involve the careful crafting of prompts, employing techniques such as persuasion (Zeng et al., 2024b), nested scene construction (Li et al., 2023), and persona modulation (Shah et al., 2023). Andriushchenko and Flammarion (2024) introduce a method that applies past-tense transformations to OpenAI’s recent o1 reasoning models, revealing their lack of robustness against subtle linguistic shifts. Ying et al. (2025b) propose attack prompts that combine common jailbreak strategies—such as scenario injection, affirmative prefixes, and indirect instructions—with safety-sensitive queries to probe model vulnerabilities. Their findings indicate that reasoning models like DeepSeek-R1 and OpenAI’s o1 are particularly susceptible to such attacks, as their explicit CoT reasoning renders them more exploitable than standard LLMs.
Multi-turn Jailbreak.
Performing jailbreak attacks in a single query can be challenging, but multi-turn conversations or sequential prompts may incrementally guide models toward generating restricted content Russinovich et al. (2024); Sun et al. (2024). Multi-turn attacks are particularly relevant to reasoning-capable models as these models possess sophisticated logical processing that can be exploited through extended dialogues. Ying et al. (2025a) propose Reasoning-Augmented Conversation (RACE), which reformulates harmful queries into benign reasoning tasks and gradually exploits the model’s inference capabilities to compromise safety alignment, achieving success rates up to 96%. Ren et al. (2024) introduce ActorAttack, a framework that constructs semantically linked conversational sequences that appear harmless individually but collectively lead to harmful outputs, successfully targeting even advanced models like o1. Li et al. (2024) further show that multi-turn human jailbreaks significantly outperform automated single-turn attacks, leveraging the model’s ability to maintain context and be incrementally steered toward unsafe behaviors.
Reasoning Exploitation Jailbreak.
LRMs possess advanced reasoning capabilities that, while enhancing their utility, introduce unique vulnerabilities that can be exploited through reasoning-based jailbreak attacks. Unlike traditional LLMs, these models explicitly expose their CoT reasoning processes, creating new attack surfaces.  Yao et al. (2025) introduce Mousetrap, a framework that leverages chaos mappings to create iterative reasoning chains that gradually lead LRMs into harmful outputs. By embedding one-to-one mappings into the reasoning process, Mousetrap effectively traps models like OpenAI’s o1-mini and Claude-sonnet with success rates of up to 98%. Kuo et al. (2025) propose Hijacking Chain-of-Thought (H-CoT), which manipulates the reasoning process by injecting execution-phase thoughts that bypass safety checks entirely. Their approach exploits LRMs’ tendency to prioritize problem-solving over safety considerations, causing rejection rates to plummet from 98% to below 2% across models like OpenAI o1/o3 and DeepSeek-R1. Both approaches demonstrate that the very reasoning mechanisms designed to enhance LRMs’ capabilities can become their most significant security weaknesses when strategically manipulated.
5 Defenses for LRMs
To mitigate safety risks and defend against attacks on LRMs, various defense strategies have been proposed in recent research. We categorize these approaches into three main types: Safety Alignment (Section 5.1), Inference-Time Defenses (Section 5.2), and Guard Models (Section 5.3).
5.1 Safety Alignment of LRMs
Similar to LLMs and VLMs, LRMs are required to align with humans’ values and expectations. The 3H principle (Askell et al., 2021) (Helpful, Honest, and Harmless) provides a foundational guideline for constraining model behaviors.
The existing safety alignment pipelines and techniques developed for LLMs (Shen et al., 2023) and VLMs (Ye et al., 2025) can be readily adapted to LRMs, as they share similar architectures and natural language generation behaviors. For example, the alignment process for LLMs typically starts with collecting high-quality, value-aligned data (Ethayarajh et al., 2022), either from existing benchmarks (Bach et al., 2022; Wang et al., 2022c), LLM-generated instructions (Wang et al., 2022b), or by filtering unsafe content (Welbl et al., 2021; Wang et al., 2022a). During training, common techniques include supervised fine-tuning (SFT) (Wu et al., 2021), reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), and direct preference optimization (DPO) (Rafailov et al., 2024). In the domain of VLMs, safety alignment has been achieved through various approaches. For example, Liu et al. (2024) introduce additional safety modules during training to enhance model alignment. Moreover, methods such as ADPO (Weng et al., 2025), Safe RLHF-V (Ji et al., 2025), and GRPO-based methods (Li et al., 2025a) improve safety via DPO (Rafailov et al., 2024), RLHF (Ouyang et al., 2022), and GRPO (DeepSeek-AI et al., 2025), respectively. Additionally, open-source datasets and benchmarks (Zhang et al., 2024; Ji et al., 2025) have played a crucial role in providing high-quality alignment data for safety evaluation.
Although effective, the previous alignment methods for LLMs and VLMs may overlook the reasoning process of LRMs, leading to unsatisfactory alignment performance. To mitigate this challenge, various works focus on different aspects, including safe CoT data curation, SFT-based safety alignment on reasoning, and RL-based safety alignment on reasoning.
Safe CoT Data Curation.
First, Wang et al. (2025) build a 1k-scale safety dataset named STAR-1 specifically designed for LRMs. Another safety training data in CoT style named SafeChain (Jiang et al., 2025) is introduced to enhance the safety of LRMs. In addition, Zhang et al. (2025b) construct a dataset consisting of 15k safety-aware reasoning trajectories, generated by DeepSeek-R1, with explicit instructions designed to promote expected refusal behavior.
SFT-based Safety Alignment on Reasoning.
Based on the curated safe CoT data, researchers further conduct SFT to improve safety. For example, Jiang et al. (2025) train two LRMs with the SafeChain dataset, demonstrating that it not only enhances model safety but also preserves reasoning performance. Besides, RealSafe-R1 (Zhang et al., 2025b) is developed to make LRMs safer by training DeepSeek-R1 distilled models on the 15k safety-aware reasoning trajectories.
RL-based Safety Alignment on Reasoning.
In addition to SFT, various further post-training techniques for safety are proposed based on reinforcement learning (RL). For example, deliberative alignment (Guan et al., 2024) is proposed to teach models safety specifications directly and train them to reason over these guidelines before generating responses explicitly via reinforcement learning. In addition, STAIR (Zhang et al., 2025c) utilizes Monte Carlo tree search and DPO (Rafailov et al., 2024) to integrate safety alignment with introspective reasoning. Meanwhile, a SaRO (Mou et al., 2025) is proposed to incorporate safety-policy-driven reasoning into the alignment process. Besides, R2D (Zhu et al., 2025a) is present to unlock the safety-aware reasoning mechanism to defense against jailbreak attacks with the proposed contrastive pivot optimization (CPO).
However, safety alignment brings the safety alignment tax (Lin et al., 2023a), compromising the fundamental capabilities of LRMs like reasoning capability (Huang et al., 2025). To mitigate this issue, researchers explore alternative defense techniques that do not require direct modifications to the victim models.
5.2 Inference-time Defenses for LRMs
To circumvent the safety alignment tax (Lin et al., 2023a; Huang et al., 2025), one line of work focuses on applying defenses at inference time. The insights from previous inference-time defenses for LLMs (Cheng et al., 2023; Lu et al., 2023) and VLMs (Wang et al., 2024a; Ghosal et al., 2024; Ding et al., 2024; Liu et al., 2025a), such as safe system prompting, few-shot safe demonstrations, and safe decoding, can be naturally borrowed to LRMs, as the token generation mechanism is similar across these models.
However, the reasoning process in LRMs brings new challenges and opportunities for inference-time defenses. Therefore, various inference-time techniques like inference-time scaling on reasoning and safe decoding for reasoning are proposed to ensure the safety of reasoning in LRMs.
Inference-time Scaling on Reasoning.
Zaremba et al. (2025) demonstrate that the inference-time scaling on reasoning improves the safety and adversarial robustness of LRMs. Future work could explore dynamic scaling strategies tailored to input complexity, or integrate adaptive reasoning depth control to balance efficiency and safety performance (Liu et al., 2025c) during inference.
Safe Decoding for Reasoning.
Jiang et al. (2025) propose three decoding strategies, including ZeroThink, LessThink, and MoreThink, to verify model safety during reasoning. Making the reasoning safer at inference time could be a promising future direction, by verifying intermediate steps, filtering unsafe trajectories, or integrating reasoning-aware guard mechanisms during decoding.
5.3 Guard Models for LRMs
Another line of work without direct modification to the victim model focuses on building guard models for the victim model. The previous inference-time defenses still focus on the safer inference of the victim models themselves. Differently, guard models aim to moderate the input and output of the victim models without training the victim models or modifying the inference strategies of the victim models. The existing guard models for LLMs (Inan et al., 2023) or VLMs (Chi et al., 2024b) can also safeguard the LRMs since they share similar input and output formats. In addition, the reasoning-based guard models (Liu et al., 2025b) can better moderate the reasoning process of LRMs via guiding the guard models to deliberatively reason before making moderation decisions. We category existing guard models into two classes, including classifier-based guard models and reasoning-based guard models.
Classifier-based Guard Models.
The LLM guard models, including ToxicChat-T5 (Lin et al., 2023b), ToxDectRoberta (Zhou, 2020), LaGoNN (Bates and Gurevych, 2023), the LLaMA Guard series (Inan et al., 2023; Dubey et al., 2024), Aegis Guard series (Ghosh et al., 2024a, b), WildGuard (Han et al., 2024), ShieldGemma (Zeng et al., 2024a), are typically based on open-sourced LLMs and fine-tuned on the red-teaming data. In the VLM domain, for example, LLaVAGuard (Helff et al., 2024) is built to conduct large-scale dataset annotation and moderate the text-image models. In addition, VLMGuard (Du et al., 2024) is proposed to conduct malicious image-text prompt detection by leveraging the unlabeled user prompts. Moreover, LLaMA Guard 3-Vision (Chi et al., 2024a) is developed to moderate both the image-text input and text output of VLMs via SFT. To improve the generalization ability, (Ji et al., 2025) presents Beaver-Guard-V by training a reward model and then applying reinforcement learning. Although effective, they are typically classifier-based guard models, limiting their abilities in moderate reasoning data. To mitigate this problem, the reasoning-based guard models (Liu et al., 2025b) are proposed to enhance the reasoning ability of guard models.
Reasoning-based Guard Models.
Through the proposed reasoning SFT and hard sample DPO, GuardReasoner (Liu et al., 2025b) is proposed to guide the guard model to deliberatively reason before making moderation decisions, improving performance, generalization ability, and explainability. Similarly, ThinkGuard (Wen et al., 2025) is developed via the proposed critique-augmented fine-tuning. X-Guard (Upadhayay et al., 2025) extends the reasoning-based guard model to the multi-lingual scenario.
6 Future Directions
Beyond the detailed analysis of risks, attacks, and defenses presented in previous sections, this paper also identifies future directions that researchers should prioritize to enhance the safety of LRMs: (1) Standardized Evaluation Benchmarks. New benchmarks should focus on reasoning-specific vulnerabilities, as the research community currently lacks standardized evaluation frameworks to comprehensively test both the safety and robustness of LRMs’ multi-step reasoning processes. (2) Domain-Specific Evaluation Frameworks. Evaluation suites for healthcare, finance, and law must include curated case studies and targeted adversarial tests. Expert review ensures LRMs meet each domain’s accuracy and ethical requirements. (3) Human-in-the-Loop Alignment and Interpretability. Interactive tools should let experts inspect and refine reasoning traces. Iterative feedback can align LRMs with stakeholder values and correct biases efficiently.
7 Conclusion
This survey has comprehensively examined the emerging safety challenges posed by LRMs. We’ve identified unique vulnerabilities in these models that extend beyond traditional LLMs, mapping out the landscape of safety risks, adversarial attack vectors, and defense strategies. By organizing these elements into a detailed taxonomy, this work aims to facilitate future research that can enhance the security and reliability of these increasingly powerful AI systems while preserving their remarkable reasoning capabilities.
Limitations
This survey has inherent limitations due to the rapidly evolving nature of LRMs. Since the emergence of OpenAI’s o1 series, DeepSeek-R1, and other advanced reasoning models is relatively recent, our taxonomy and findings may become outdated as new research continuously emerges. While we have endeavored to provide a comprehensive overview of safety challenges, attacks, and defenses, we acknowledge that some aspects may require revision as the field matures. Additionally, our reliance on published academic literature may not fully capture proprietary research being conducted within companies developing these models, potentially creating gaps in understanding industry-specific safety measures.